# Hardware-inspired Continuous Variables Quantum Optical Neural Networks

### v1

@@ 1 Introduction
Continuous-variable (CV) quantum optics serves as a natural environment to implement analogs of classical neural networks by exploiting the infinite-dimensional Hilbert space of optical modes. In the scheme of CV quantum optical neural networks (QONN), Gaussian unitaries realized with displacements, squeezers, beam splitters and phase shifters lay the role of affine transformations on the optical quadratures—continuous degrees of freedom of photons—, while non-Gaussian resources supply the essential activation—nonlinear—function and universality for CV quantum information processing lloyd1999quantum.
Early CV-QONN proposals realized neurons by alternating Gaussian layers with explicit non-Gaussian unitaries as Kerr gate, thereby enforcing nonlinearity and universality of CV quantum computing lloydcvqnn; steinbrecher2019quantum. A complementary direction pursued measurement-induced nonlinearity: Siopsis et al. used only Gaussian unitaries with ancillas and a homodyne repeat-until-success protocol to impart non-Gaussian kicks to the data modes siopsis2023exprealcvqnn, while Chabaud and Sohbi studied adaptive linear optics where each layer applies an interferometer followed by photon counting with feed-forward chabaud2021quantum.
More recently, programmable nonlinear meshes have been introduced by embedding tunable Kerr-like elements directly into interferometer networks chernykh2024programmable, and particle-number preserving photonic convolutional NNs have employed state injection of non-Gaussian ancilla at designated layers monbroussou2025photonic.
@@
The practical development of a CV-QONN hinges on balancing experimental overhead with the non-Gaussian resources needed for universality. Non-Gaussian unitaries such as the Kerr gate, while universal in principle, are widely regarded as hardware-demanding at optical frequencies because the third-order susceptibility $\chi^{(3)}$ yields very small per-photon phase shifts, making high-fidelity deterministic implementations challenging shapiro2006single. Measurement-induced cubic operations are generally possible, but they require tailored non-Gaussian ancilla states and high-performance feed-forward, which remains experimentally demanding marek2011deterministic. However, the state-engineering literature established heralded photon subtraction as a practical non-Gaussian resource: tapping a weak fraction of squeezed light and conditioning on a photon detection event reliably produces non-Gaussian states such as cat states and odd-photon superpositions ourjoumtsev2006science; neergaard2006prl; wakui2007oe; olivares2005ips. Recent approaches, such as generalized photon subtraction with photon-number-resolving (PNR) detection, further improve cat-state generation rates and operating regimes takase2021pra.
Inspired by the variational CV circuit of Stornati et al. stornati2024variational, this work presents the CV-QONN architecture from Fig. 1 where, in a quantum system of several optical modes, initial coherent states encode the input, parameterized Gaussian optics—interferometers $\hat{U}(\bm{\phi},\bm{\eta},\bm{\varphi})$, squeezers $\hat{S}(\bm{r},\bm{\vartheta})$, and displacements $\hat{D}(\bm{\delta})$—are assembled in a general Gaussian operator $\hat{G}(\bm{\theta})$ and realize the affine part of each layer, photon subtraction $\hat{a}$ provides the nonlinearity, and homodyne detection of position quadrature $\langle\hat{x}\rangle$ implements the output readout. The entire stack of quantum operators maps directly onto hardware-friendly components: local oscillators, beam splitters, phase shifters, optical parametric amplifiers for single-mode squeezers, single-photon detectors and photodetectors.
A key feature of the introduced design is that, for quantum states built with Gaussian operators and a finite number of annihilation/creation operators, expectation values of observables admit exact analytic expressions via Bogoliubov transformations and Wick–Isserlis expansion in terms of the first and second statistical moments of the underlying Gaussian state lloyd2012gaussian; ulysse2021classical. This allows closed-form expression derivations on genuinely non-Gaussian states generated by photon subtraction/addition operations, which enable gradient-based training and avoids Fock-space cutoffs that are common in classical simulations of CV quantum systems.
@@ 2 Overview
The main theoretical result is the derivation of the closed-form expression in Eq. (17)—generalized for multi-mode subtractions on a general Gaussian state in Eq. (27)—that establishes the emergence of non-polynomial, tunable ridge activations induced by photon subtraction on squeezed coherent states. This analytic mapping reveals that each photon-subtracted mode corresponds to an effective quantum-optical neuron. Consequently, a single QONN layer with linear readout satisfies the Universal Approximation Theorem (UAT) on compact sets cybenko1989approximation; leshno1993multilayer. When several subtractions occur over correlated squeezed modes, additional nonlinear cross-terms appear, enhancing the expressivity of the network.
From a practical viewpoint, all required photonic components—beam splitters, phase shifters, squeezers, heralded photon-subtraction modules, and homodyne detection—are experimentally accessible with current technology. Limitations such as the available squeezing strength ($|r|\leq 1.7$) vahlbruch2024squeezinglims are mitigated by the tunable affine mapping stage, ensuring the inputs lie within the effective nonlinear regime.
@@
  • Classification: nonlinear decision boundaries such as in the “moons” and “circles” datasets are correctly learned by single- and two-neuron QONNs, achieving perfect accuracy. A multi-class MNIST classification task with continuous encodings confirms scalability to realistic data.
@@ 3.1 Feed-forward neural networks
where $W\in\mathbb{R}^{K\times A}$ is the weight matrix that linearly transforms $\bm{\alpha}$ along with the bias vector $\bm{d}\in\mathbb{R}^{K}$, while $V\in\mathbb{R}^{M\times K}$ and $\bm{c}\in\mathbb{R}^{M}$ provide the linear readout of the layer, yielding the outputs $\{\mathcal{L}_{j}(\bm{\alpha})\}_{j=1}^{M}$. Each term $\Phi(W_{k}\bm{\alpha}+\bm{d}_{k})$ depends on the input only through the one-dimensional projection $W_{k}\bm{\alpha}$, which consequently corresponds to a ridge function. The amount of ridge functions $K$ in the layer determines the number of neurons and therefore the width of the layer. This enables the Universal Approximation Theorem (UAT) when $\Phi$ is non-polynomial, meaning that a single-layer NN can approximate any continuous function with a decreasing error when growing the width of the layer cybenko1989approximation; leshno1993multilayer.
@@ 3.2 CV quantum optics
must hold to preserve the canonical commutation relations (CCR) in the system serafini2017quantum. For the purpose of this work, the characterization of quantum optical systems will be carried out in Fock space with $\bm{\hat{b}}$.
@@
On the one hand, Gaussian quantum states are referred to systems governed by, at most, quadratic Hamiltonians and whose phase-space of their quadratures can be represented by a Gaussian distribution serafini2017quantum.
@@
implying that classical systems can efficiently simulate Gaussian quantum states lloyd2012gaussian; serafini2017quantum.
Gaussian operators correspond to unitary operators in Fock space and affine transformations in phase-space. Any Gaussian transformation acting on a Gaussian state preserves the system’s Gaussianity, which also implies that the product of Gaussian operators yields another Gaussian operator lloyd2012gaussian; ferraro2005gaussian.
@@
$\hat{D}_{\bm{\delta}}\rho_{G}\hat{D}_{\bm{\delta}}^{\dagger}:\begin{cases}\bm{\bar{b}}\rightarrow\bm{\bar{b}}+\bm{d}\\ \Sigma\rightarrow\Sigma\end{cases}\text{with }\bm{d}=(\bm{\delta},\bm{\delta}^{*})^{\mathsf{T}}.$ (7)
@@
$\begin{split}&\hat{Q}\rho_{G}\hat{Q}^{\dagger}:\begin{cases}\bm{\bar{b}}\rightarrow B\bm{\bar{b}}\\ \Sigma\rightarrow B\Sigma B^{\dagger}\end{cases}\text{where}\\ &B=\begin{pmatrix}U&V\\ V^{*}&U^{*}\end{pmatrix}\text{ and }\begin{cases}U=U_{2}\cosh R\ U_{1}^{\dagger}\\ V=-U_{2}e^{i2\varTheta}\sinh R\ U_{1}^{\mathsf{T}}\end{cases},\end{split}$ (8)
@@
Thus, the most general $N$-mode Gaussian operator $\hat{G}$ is defined by joining both linear and quadratic Gaussians lloyd2012gaussian in the sequence
@@
On the other hand, non-Gaussian quantum states cannot be identified only by their first two statistical moments, and, in the general case, their classical tractability goes beyond efficient Gaussian simulation. They are a necessary resource to achieve universal quantum computation in CV systems walschaers2021non.
Non-Gaussian operators, under some conditions, tend to introduce statistical correlations of order higher than two and can inject nonlinear phase-space dynamics into the quantum state they act on lloydcvqnn, analogous to activation functions in the classical NN framework.
In contrast to physically demanding non-Gaussian operators such as Kerr or cubic-phase gates, this work uses a hardware-accessible method to generate non-Gaussianity in a quantum optical system via the non-unitary particle-annihilation—photon subtraction—operator $\hat{a}$ biagi2025.
For a non-Gaussian state built as a polynomial of ladder operators acting on some pure Gaussian state $\rho_{G}$, Wick–Isserlis theorem decomposes any expectation value of such state construction into a sum of products of linear and quadratic ladder expectations, which can be collected from the means and covariances of $\rho_{G}$ stornati2024variational; nongauss2023emergent. The general Wick–Isserlis expansion for non-zero means nosal2022higher applied to this framework is defined as
@@ 4.1 Gaussianity as affine mapping
Let $\bm{\alpha}\in\mathbb{R}^{M}$ be the $M$-dimensional input vector to be uploaded, $\hat{G}$ from Eq. (9) be the $N$-mode general Gaussian unitary, and Eqs. (7) and (LABEL:quad-gaussian-op) be the actions of Gaussian transformations over an arbitrary $N$-mode Gaussian state. The $N$-mode quantum system starts in the vacuum—Gaussian—state $\rho_{\text{v}}$ defined by the null vector of means $\bm{\bar{b}}_{\text{v}}$ and the normalized identity covariance matrix $\Sigma_{\text{v}}$ lloyd2012gaussian like
@@
$\rho_{\bm{\alpha}}=\hat{D}_{\bm{\alpha}}\rho_{\text{v}}\hat{D}_{\bm{\alpha}}^{\dagger}\equiv\begin{cases}\bm{\bar{b}}_{\bm{\alpha}}=(\bm{\alpha},\bm{0}^{N-M},\bm{\alpha}^{*},\bm{0}^{N-M})\\ \Sigma_{\bm{\alpha}}=\Sigma_{\text{v}}\end{cases},$ (13)
@@ 4.2 Subtraction as nonlinear mapping
$\rho_{r\alpha}\equiv\begin{cases}\bm{\bar{b}}=(e^{r}\alpha,e^{r}\alpha^{*})\\ \Sigma=\begin{pmatrix}\cosh^{2}r&\frac{1}{2}\sinh 2r\\ \frac{1}{2}\sinh 2r&\sinh^{2}r\end{pmatrix}\end{cases}$, (16)
@@
These features make Eq. (17) a suitable activation function for the QONN whose bump gain and form can be controlled by $r$, providing a complete family of nonlinear mappings of the input $\alpha$ in the $\hat{x}$ quadrature as exemplified in Fig. 2, and enabling the UAT cybenko1989approximation; leshno1993multilayer.
@@ 4.3 CV-QONN layer closed form
This closed-form establishes the evidence that photon subtractions in distinct optical modes of the system give rise to a linear combination of single-ridge activations—effectively one quantum-optical neuron per photon-subtracted mode—so a single-layer QONN with linear readout achieves the UAT on compact sets cybenko1989approximation; leshno1993multilayer. When two or more photon-subtracted modes are correlated by non-zero off-diagonal covariances induced by the passive mixer $U_{2}$ after the squeezing, additional nonlinear cross terms appear enriching the quantum neuron operation, thus increasing the QONN expressivity and potentially reducing the width $|K|$ needed to reach a target accuracy.
@@ 4.4 Practical implications
With respect to hardware limitations, all the quantum optical elements employed are physically available except the strength of the squeezing operator, which is currently restricted to certain ranges when it comes to its realization in the laboratory schnabel2017. Recent works as vahlbruch2024squeezinglims reported a ceiling of $15$ dB $\approx|r|=1.73$ for the squeezing factor. Although the optimal nonlinear range for $\alpha$ seen in Eq. (20) is constrained by the squeezing strength $r$ and by the lower-bound of nonlinear contribution $\varepsilon$, the trainable affine mapping from Eq. (14) would always bring the uploaded inputs $\bm{\alpha}$ to a convenient regime, where the nonlinearities of the subtractions have a non-trivial activation effect on the optical quadratures, avoiding the necessity of setting an explicit range for $\bm{\alpha}$.
Lastly, the physical implementation of the QONN is intrinsically probabilistic and conditioned on the success probability of the heralded photon-subtraction events. The hardware realization of the annihilation operator in a given signal mode consists in mixing the target mode $\hat{a}_{k}$ with an ancillary vacuum mode $\hat{b}$ on a weak tap beam splitter of signal reflectivity $\eta\ll 1$, followed by single-photon detection on the ancilla arm ourjoumtsev2006science; dakna1997generating; wenger2004nonGaussian; garciapatron2004photon. The corresponding two-mode beam splitter unitary acts as
@@
consistent with multi-mode analysis of engineered addition/subtraction sequences fiurasek2009engineering. In practice, one may fold detector inefficiency, optical loss and mode-matching into an effective tap $\eta\to\eta_{\mathrm{eff}}$ without changing the form of the expressions above.

### v2

@@ 1 Introduction
Continuous-variable (CV) quantum optics serves as a natural environment to implement analogs of classical neural networks by exploiting the infinite-dimensional Hilbert space of optical modes. In the scheme of CV quantum optical neural networks (QONN), Gaussian unitaries realized with displacements, squeezers, beam splitters and phase shifters lay the role of affine transformations on the optical quadratures—continuous degrees of freedom of photons—, while non-Gaussian resources supply the essential activation—nonlinear—function and universality for CV quantum information processing Lloyd and Braunstein (1999).
Early CV-QONN proposals realized neurons by alternating Gaussian layers with explicit non-Gaussian unitaries as Kerr gate, thereby enforcing nonlinearity and universality of CV quantum computing Killoran et al. (2019a); Steinbrecher et al. (2019). A complementary direction pursued measurement-induced nonlinearity: Siopsis et al. used only Gaussian unitaries with ancillas and a homodyne repeat-until-success protocol to impart non-Gaussian kicks to the data modes Bangar et al. (2023), while Chabaud et al. studied adaptive linear optics where each layer applies an interferometer followed by photon counting with feed-forward Chabaud et al. (2021b).
More recently, programmable nonlinear meshes have been introduced by embedding tunable Kerr-like elements directly into interferometer networks Chernykh et al. (2024), and particle-number preserving photonic convolutional NNs have employed state injection of non-Gaussian ancilla at designated layers Monbroussou et al. (2025).
@@
The practical development of a CV-QONN hinges on balancing experimental overhead with the non-Gaussian resources needed for universality. Non-Gaussian unitaries such as the Kerr gate, while universal in principle, are widely regarded as hardware-demanding at optical frequencies because the third-order susceptibility $\chi^{(3)}$ yields very small per-photon phase shifts, making high-fidelity deterministic implementations challenging Shapiro (2006). Measurement-induced cubic operations are generally possible, but they require tailored non-Gaussian ancilla states and high-performance feed-forward, which remains experimentally demanding Marek et al. (2011). However, the state-engineering literature established heralded photon subtraction as a practical non-Gaussian resource: tapping a weak fraction of squeezed light and conditioning on a photon detection event reliably produces non-Gaussian states such as cat states and odd-photon superpositions Ourjoumtsev et al. (2006); Neergaard-Nielsen et al. (2006); Wakui et al. (2007); Olivares and Paris (2005). Recent approaches, such as generalized photon subtraction with photon-number-resolving (PNR) detection, further improve cat-state generation rates and operating regimes Takase et al. (2021).
Inspired by the variational CV circuit of Stornati et al. Stornati et al. (2024), this work presents the CV-QONN architecture from Fig. 1 where, in a quantum system of several optical modes, initial coherent states encode the input, parameterized Gaussian optics—interferometers $\hat{U}(\bm{\phi},\bm{\eta},\bm{\varphi})$, squeezers $\hat{S}(\bm{r},\bm{\vartheta})$, and displacements $\hat{D}(\bm{\delta})$—are assembled in a general Gaussian operator $\hat{G}(\bm{\theta})$ and realize the affine part of each layer, photon subtraction $\hat{a}$ provides the nonlinearity, and homodyne detection of position quadrature $\langle\hat{x}\rangle$ implements the output readout. The entire stack of quantum operators maps directly onto hardware-friendly components: local oscillators, beam splitters, phase shifters, optical parametric amplifiers for single-mode squeezers, single-photon detectors and photodetectors.
A key feature of the introduced design is that, for quantum states built with Gaussian operators and a finite number of annihilation/creation operators, expectation values of observables admit exact analytic expressions via Bogoliubov transformations and Wick–Isserlis expansion in terms of the first and second statistical moments of the underlying Gaussian state Weedbrook et al. (2012); Chabaud et al. (2021a). This allows closed-form expression derivations on genuinely non-Gaussian states generated by photon subtraction/addition operations, which enable gradient-based training and avoids Fock-space cutoffs that are common in classical simulations of CV quantum systems.
@@ 2 Overview
The main theoretical result is the derivation of the closed-form expression in Eq. (17)—generalized for multi-mode subtractions on a general Gaussian state in Eq. (27)—that establishes the emergence of non-polynomial, tunable ridge activations induced by photon subtraction on squeezed coherent states. This analytic mapping reveals that each photon-subtracted mode corresponds to an effective quantum-optical neuron. Consequently, a single QONN layer with linear readout satisfies the Universal Approximation Theorem (UAT) on compact sets Cybenko (1989); Leshno et al. (1993). When several subtractions occur over correlated squeezed modes, additional nonlinear cross-terms appear, enhancing the expressivity of the network.
From a practical viewpoint, all required photonic components—beam splitters, phase shifters, squeezers, heralded photon-subtraction modules, and homodyne detection—are experimentally accessible with current technology. Limitations such as the available squeezing strength ($|r|\leq 1.7$) Vahlbruch et al. (2024) are mitigated by the tunable affine mapping stage, ensuring the inputs lie within the effective nonlinear regime.
@@
  • Classification: nonlinear decision boundaries such as in the “moons” and “circles” datasets are correctly learned by single-neuron QONNs with two and three modes, achieving perfect accuracy. A multi-class MNIST classification task with continuous encodings confirms scalability to realistic data.
@@ 3.1 Feed-forward neural networks
where $W\in\mathbb{R}^{K\times A}$ is the weight matrix that linearly transforms $\bm{\alpha}$ along with the bias vector $\bm{d}\in\mathbb{R}^{K}$, while $V\in\mathbb{R}^{M\times K}$ and $\bm{c}\in\mathbb{R}^{M}$ provide the linear readout of the layer, yielding the outputs $\{\mathcal{L}_{j}(\bm{\alpha})\}_{j=1}^{M}$. Each term $\Phi(W_{k}\bm{\alpha}+\bm{d}_{k})$ depends on the input only through the one-dimensional projection $W_{k}\bm{\alpha}$, which consequently corresponds to a ridge function. The amount of ridge functions $K$ in the layer determines the number of neurons and therefore the width of the layer. This enables the Universal Approximation Theorem (UAT) when $\Phi$ is non-polynomial, meaning that a single-layer NN can approximate any continuous function with a decreasing error when growing the width of the layer Cybenko (1989); Leshno et al. (1993).
@@ 3.2 CV quantum optics
must hold to preserve the canonical commutation relations (CCR) in the system Serafini (2017). For the purpose of this work, the characterization of quantum optical systems will be carried out in Fock space with $\bm{\hat{b}}$.
@@
On the one hand, Gaussian quantum states are referred to systems governed by, at most, quadratic Hamiltonians and whose phase-space of their quadratures can be represented by a Gaussian distribution Serafini (2017).
@@
implying that classical systems can efficiently simulate Gaussian quantum states Weedbrook et al. (2012); Serafini (2017).
Gaussian operators correspond to unitary operators in Fock space and affine transformations in phase-space. Any Gaussian transformation acting on a Gaussian state preserves the system’s Gaussianity, which also implies that the product of Gaussian operators yields another Gaussian operator Weedbrook et al. (2012); Ferraro et al. (2005).
@@
$\hat{D}_{\bm{\delta}}\rho_{G}\hat{D}_{\bm{\delta}}^{\dagger}:\begin{cases}\bm{\bar{b}}\rightarrow\bm{\bar{b}}+\bm{d}\\ \Sigma\rightarrow\Sigma\end{cases}\text{with }\bm{d}=(\bm{\delta},\bm{\delta}^{\ast})^{\mathsf{T}}.$ (7)
@@
$\begin{split}&\hat{Q}\rho_{G}\hat{Q}^{\dagger}:\begin{cases}\bm{\bar{b}}\rightarrow B\bm{\bar{b}}\\ \Sigma\rightarrow B\Sigma B^{\dagger}\end{cases}\text{where}\\ &B=\begin{pmatrix}U&V\\ V^{\ast}&U^{\ast}\end{pmatrix}\text{ and }\begin{cases}U=U_{2}\cosh R\ U_{1}^{\dagger}\\ V=-U_{2}e^{i2\varTheta}\sinh R\ U_{1}^{\mathsf{T}}\end{cases},\end{split}$ (8)
@@
Thus, the most general $N$-mode Gaussian operator $\hat{G}$ is defined by joining both linear and quadratic Gaussians Weedbrook et al. (2012) in the sequence
@@
On the other hand, non-Gaussian quantum states cannot be identified only by their first two statistical moments, and, in the general case, their classical tractability goes beyond efficient Gaussian simulation. They are a necessary resource to achieve universal quantum computation in CV systems Walschaers (2021).
Non-Gaussian operators, under some conditions, tend to introduce statistical correlations of order higher than two and can inject nonlinear phase-space dynamics into the quantum state they act on Killoran et al. (2019a), analogous to activation functions in the classical NN framework.
In contrast to physically demanding non-Gaussian operators such as Kerr or cubic-phase gates, this work uses a hardware-accessible method to generate non-Gaussianity in a quantum optical system via the non-unitary particle-annihilation—photon subtraction—operator $\hat{a}$ Biagi et al. (2025).
For a non-Gaussian state built as a polynomial of ladder operators acting on some pure Gaussian state $\rho_{G}$, Wick–Isserlis theorem decomposes any expectation value of such state construction into a sum of products of linear and quadratic ladder expectations, which can be collected from the means and covariances of $\rho_{G}$ Stornati et al. (2024); Walschaers et al. (2023). The general Wick–Isserlis expansion for non-zero means Nosal and Teretenkov (2022) applied to this framework is defined as
@@ 4.1 Gaussianity as affine mapping
Let $\bm{\alpha}\in\mathbb{R}^{M}$ be the $M$-dimensional input vector to be uploaded, $\hat{G}$ from Eq. (9) be the $N$-mode general Gaussian unitary, and Eqs. (7) and (LABEL:quad-gaussian-op) be the actions of Gaussian transformations over an arbitrary $N$-mode Gaussian state. The $N$-mode quantum system starts in the vacuum—Gaussian—state $\rho_{\text{v}}$ defined by the null vector of means $\bm{\bar{b}}_{\text{v}}$ and the normalized identity covariance matrix $\Sigma_{\text{v}}$ Weedbrook et al. (2012) like
@@
$\rho_{\bm{\alpha}}=\hat{D}_{\bm{\alpha}}\rho_{\text{v}}\hat{D}_{\bm{\alpha}}^{\dagger}\equiv\begin{cases}\bm{\bar{b}}_{\bm{\alpha}}=(\bm{\alpha},\bm{0}^{N-M},\bm{\alpha}^{\ast},\bm{0}^{N-M})\\ \Sigma_{\bm{\alpha}}=\Sigma_{\text{v}}\end{cases},$ (13)
@@ 4.2 Subtraction as nonlinear mapping
$\rho_{r\alpha}\equiv\begin{cases}\bm{\bar{b}}=(e^{r}\alpha,e^{r}\alpha^{\ast})\\ \Sigma=\begin{pmatrix}\cosh^{2}r&\frac{1}{2}\sinh 2r\\ \frac{1}{2}\sinh 2r&\sinh^{2}r\end{pmatrix}\end{cases}$, (16)
@@
These features make Eq. (17) a suitable activation function for the QONN whose bump gain and form can be controlled by $r$, providing a complete family of nonlinear mappings of the input $\alpha$ in the $\hat{x}$ quadrature as exemplified in Fig. 2, and enabling the UAT Cybenko (1989); Leshno et al. (1993).
@@ 4.3 CV-QONN layer closed form
This closed-form establishes the evidence that photon subtractions in distinct optical modes of the system give rise to a linear combination of single-ridge activations—effectively one quantum-optical neuron per photon-subtracted mode—so a single-layer QONN with linear readout achieves the UAT on compact sets Cybenko (1989); Leshno et al. (1993). When two or more photon-subtracted modes are correlated by non-zero off-diagonal covariances induced by the passive mixer $U_{2}$ after the squeezing, additional nonlinear cross terms appear enriching the quantum neuron operation, thus increasing the QONN expressivity and potentially reducing the width $|K|$ needed to reach a target accuracy.
@@ 4.4 Practical implications
With respect to hardware limitations, all the quantum optical elements employed are physically available except the strength of the squeezing operator, which is currently restricted to certain ranges when it comes to its realization in the laboratory Schnabel (2017). Recent works as Vahlbruch et al. (2024) reported a ceiling of $15$ dB $\approx|r|=1.73$ for the squeezing factor. Although the optimal nonlinear range for $\alpha$ seen in Eq. (20) is constrained by the squeezing strength $r$ and by the lower-bound of nonlinear contribution $\varepsilon$, the trainable affine mapping from Eq. (14) would always bring the uploaded inputs $\bm{\alpha}$ to a convenient regime, where the nonlinearities of the subtractions have a non-trivial activation effect on the optical quadratures, avoiding the necessity of setting an explicit range for $\bm{\alpha}$.
Lastly, the physical implementation of the QONN is intrinsically probabilistic and conditioned on the success probability of the heralded photon-subtraction events. The hardware realization of the annihilation operator in a given signal mode consists in mixing the target mode $\hat{a}_{k}$ with an ancillary vacuum mode $\hat{b}$ on a weak tap beam splitter of signal reflectivity $\eta\ll 1$, followed by single-photon detection on the ancilla arm Ourjoumtsev et al. (2006); Dakna et al. (1997); Wenger et al. (2004); García-Patrón et al. (2004). The corresponding two-mode beam splitter unitary acts as
@@
consistent with multi-mode analysis of engineered addition/subtraction sequences Fiurášek (2009). In practice, one may fold detector inefficiency, optical loss and mode-matching into an effective tap $\eta\to\eta_{\mathrm{eff}}$ without changing the form of the expressions above.